OpenStudy (clamin):

8. The standard deviations for the blue and yellow teams are shown below. Complete the chart by calculating the variance,   , and the standard deviation,  , for the red team. Show your work. (2 points) Red Yellow Blue Standard deviation 1.37 2.41

OpenStudy (clamin):

Red 47.5 42 48.5 44.5 47 41 48 43 49 Yellow 47.5 45.5 48.5 46.5 47 45 48 46 49 Blue 48.1 46.8 52 47.9 48 46 50 47 53

OpenStudy (anonymous):

I see no chart. Anyway, to find standard deviation, you use the formula:\[\sqrt{\frac{ \Sigma x ^{2} }{ n }-\Delta ^{2}} \], where delta is mean. Let us start with summation of xsquared. it is the sum of the square of each and every element:\[Sigmax ^{2}=47.5^2+42^2+48.5^2+44.5^2+47^2+41^2+48^2+43^2+49^2=18796.75\] n is the number of elements in the set, which is 9. \[\Delta=\frac{ \Sigma x }{ n }= \frac{ 47.5+42+48.5+44.5+47+41+48+43+49 }{ 9 }=\frac{ 821 }{ 18 }\]. make the substitutions and tell me what you get.

OpenStudy (anonymous):

Variance is the square of standard deviation.